A 0.12 g sample of a diprotic acid (H2A) require 15.00 mL of 0.30 M NaOH solution to completely react with the acid. Calculate the molar mass of the acid.
1
Expert's answer
2020-06-12T07:26:15-0400

H2A+2NaOH=Na2A+2H2O

1.we find the amount of NaOH. to do this, we multiply the concentration by the volume (liters)

0.015 litres * 0.3 mol/litres=0.0045 mol

2.based on the above reaction we can find the amount of H2A. for this we divide the amount of naoh by 2

0.0045mol /2=0.00225mol

3.now we find the molar mass of H2A. To do this, we divide the mass of H2A by the amount of matter: M=m/n=0.12g/0.00225mol=53.33g/mol≈53g/mol